### v2

# From Context to EDUs: Faithful and Structured Context Compression via Elementary Discourse Unit Decomposition

Abstract
Managing extensive context remains a critical bottleneck for Large Language Models (LLMs), particularly in applications like long-document question answering and autonomous agents where lengthy inputs incur high computational costs and introduce noise. Existing compression techniques often disrupt local coherence through discrete token removal or rely on implicit latent encoding that suffers from positional bias and incompatibility with closed-source APIs. To address these limitations, we introduce the EDU-based Context Compressor, a novel explicit compression framework designed to preserve both global structure and fine-grained details. Our approach reformulates context compression as a structure-then-select process. First, our LingoEDU transforms linear text into a structural relation tree of Elementary Discourse Units (EDUs) which are anchored strictly to source indices to eliminate hallucination. Second, a lightweight ranking module selects query-relevant sub-trees for linearization. To rigorously evaluate structural understanding, we release StructBench, a manually annotated dataset of 248 diverse documents. Empirical results demonstrate that our method achieves state-of-the-art structural prediction accuracy and significantly outperforms frontier LLMs while reducing costs. Furthermore, our structure-aware compression substantially enhances performance across downstream tasks ranging from long-context tasks to complex Deep Search scenarios.
**footnotetext: Equal contribution.footnotetext: Corresponding author. Emails: fanchao.qi@deeplang.ai
Figure 1: Overview of the EDU-based Context Compressor framework.
1 Introduction
Large Language Models (LLMs) have achieved remarkable progress in recent years (gpt4; llama3; deepseekai2025deepseekv3technicalreport; qwen3), demonstrating capabilities that approach or surpass human performance on diverse tasks (chen2021dialogsum; si-etal-2022-scl; si2023spokenwoz; yi2024survey; luo2025large; an-etal-2025-ultraif). These advances have empowered sophisticated applications such as long-context question answering (QA) (si-etal-2025-gateau), multi-document summarization (kryściński2022booksumcollectiondatasetslongform), and autonomous agents capable of performing deep research and complex reasoning (lei2025rhinoinsight; si2025goalplanjustwish). However, these applications typically require maintaining an extensive memory for context information, e.g., the used documents and agent-environment interactions. As the capabilities of LLMs continue to grow and their application scenarios expand, effectively managing such contextual memory becomes a critical bottleneck for long-context tasks. The accumulated and lengthy context not only incurs prohibitive computational costs but also introduces significant noisy information, which can overwhelm the model and degrade the final performance of LLMs (hua2025contextengineering20context; wang-etal-2025-document).
To better organize the lengthy context, context compression has emerged as a crucial technique to reduce input token length while preserving maximal semantic integrity. Explicit compression methods (jiang2023llmlingua; xu2024concise) attempt to reduce the length of context by removing tokens or sentences deemed less important, e.g., using an abstractive summarization model as a compressor to obtain the shorter context and maintain global meaning (xu2023recompimprovingretrievalaugmentedlms). However, these methods often operate on discrete tokens or rigid sentence boundaries, disrupting the local coherence of the text. Meanwhile, they typically focus on preserving the most important global information, while overlooking the original article’s structural information and fine-grained details. Conversely, implicit compression methods (mu2024learning; gecontext; liu2025contextcascadecompressionexploring) try to encode lengthy text into latent vectors to achieve higher compression ratios. However, recent studies (li2025admtreecompressinglengthycontext) show that implicit compression methods tend to have positional bias. This means they often ignore information from the beginning or middle of the context, focusing instead on the most noticeable content and overlooking less prominent details. Also, these implicit methods (cheng2025glyphscalingcontextwindows; wei2025deepseekocrcontextsopticalcompression) tend to lack flexibility, as they often require specially designed post-training processes or the use of latent vectors as new inputs. This limits the applicability of such techniques to advanced API-based models, e.g., GPT-4.1 (gpt-4.1).
In this work, we posit that an ideal compression strategy for long-context scenarios should be explicit to ensure flexibility while focusing on the structural information of the original context, thereby maintaining both global foresight and fine-grained details. Our design goal is to first transform a linear context sequence into a structural relation tree, where each node is strictly anchored to the source via coordinate pointers. Subsequently, we select the sub-tree relevant to the input query, then linearize it as the compressed context. In this way, our explicit compression method can be guided to retain not only the most salient global information, but also the structural relationships, fine-grained details, and coherent sentences, which are essential for faithful downstream reasoning. Therefore, we introduce the EDU-based Context Compressor, which consists of the LingoEDU and a lightweight ranking module. Specifically, the LingoEDU is inspired by rhetorical structure theory and built upon elementary discourse units (EDUs) (mann1988rhetorical). Unlike fixed tokens or sentences, EDUs are the minimal variable-length units that coherently convey one piece of information. LingoEDU aims to transform unstructured context sequences into structural relationship trees, where nodes represent EDUs and edges represent the existence of a discourse linkage and dependency relation between two EDUs. To efficiently obtain the structural relationship trees, we train the LingoEDU via a novel human-in-the-loop pipeline to obtain the training data followed by a supervised fine-tuning (SFT) stage. After obtaining the structural relation tree through the EDU-based decomposition module, we subsequently employ a ranking module to get the most useful sub-trees to achieve context compression. By taking the task instruction and the structural relation tree as input, we use a ranking model to identify task-relevant sub-trees and then linearize them into a compressed context. For instance, given a multi-doc QA task, the EDU-based decomposition module first maps the lengthy context into a structural relationship tree where nodes represent semantic EDUs and edges encode their original positional references. Then, the ranking model returns the nodes that are highly relevant to the user query, filtering out noisy content at the structural level. This step effectively prunes the full document tree into a query-specific subtree, retaining only the most relevant EDUs. Finally, this refined sub-tree is linearized and fed into the target LLM alongside the original query to generate the final response. This design allows our method to preserve global structure while capturing fine-grained details, and crucially, since it operates without latent representations, it ensures seamless compatibility with API-based models.
During the experiments, we aim to answer two pivotal questions: (1) Can the state-of-the-art LLMs effectively compress long contexts while preserving original structural information? and (2) Does such structure-aware compression tangibly reduce the hallucinations for downstream tasks? To answer these questions, we first propose a manually annotated benchmark comprising 248 documents across diverse formats to evaluate the abilities of LLMs to understand and describe structural information of the context. We find that even state-of-the-art models such as o3 (o3) still fail to fully understand the structural relationships within a given context, and prompt-based methods alone for these models struggle to effectively compress the context while preserving the original structural relations. Conversely, our well-trained EDU-based Context Compressor can effectively identify and preserve key structural elements while substantially reducing the context length. Meanwhile, we find that using the structured and compressed context from EDU-based Context Compressor not only reduces the input length for the model, but also filters out irrelevant and noisy tokens, thereby improving the model’s performance and reducing hallucinations on various tasks such as multi-document QA (bai2023longbench).
Our contributions are summarized as follows:
  • Novel Context Compression Framework: We introduce the EDU-based Context Compressor that leverages document structure to create concise, informative sub-trees tailored to specific queries. This approach preserves both global structure and local details for context compression. Also, it remains fully compatible with closed-source models like GPT-4.1.
  • New Benchmark: We release a manually annotated benchmark comprising 248 documents across diverse formats to enable precise evaluation of the abilities of LLMs to understand and describe structural information of the context.
  • SOTA Performance & Efficiency: Empirical results show that our method significantly outperforms advanced LLMs (e.g., o3) in understanding structural relations within the provided context and surpasses commercial APIs like Firecrawl, notably with a much lower cost.
  • Reducing Hallucinations for Long-context Tasks: We demonstrate that our proposed EDU-based Context Compressor can improve the final performance and reduce hallucinations across diverse long-context tasks, e.g., multi-document QA, summarization, and search agent scenarios. For example, our method outperforms standard baselines (e.g., +14.94% on HotpotQA within LongBench) by preserving precise evidence chains. In Deep Search tasks, it significantly enhances the performance, boosting DeepSeek-R1 by over 51.11% relatively on the HLE benchmark.
2 Methodology
We propose the EDU-based Context Compressor, a novel framework designed to achieve faithful and structural-aware context compression. As argued in the Introduction, implicit processing often suffers from positional bias and lack of transparency. Therefore, our approach adheres to an explicit compression paradigm: it transforms the linear context sequence into a Structural Relation Tree, where semantic units are strictly anchored to the source text via coordinate pointers. As illustrated in Figure 1, the framework operates as a plug-and-play module compatible with any LLM (including API-based models). It consists of two cascaded components: the LingoEDU, which parses the document into discourse-connected units, and a Ranking Module, which identifies and linearizes the most relevant sub-trees to reconstruct a high-density context.
2.1 Overall Framework
Given a long input document $\mathcal{D}$ (or a set of documents) and a user query $q$, our goal is to overcome the limitations of fixed-size chunking and latent encoding. We reformulate the context compression task as a “Structure-then-Select” process:
  1. 1.
    Phase I: Structural Decomposition. The Decomposer transforms the unstructured linear text $\mathcal{D}$ into a Structural Relation Tree $\mathcal{T}=(\mathcal{V},\mathcal{E})$.
    • Nodes ($\mathcal{V}$): represent Elementary Discourse Units (EDUs), the minimal variable-length units capable of conveying coherent semantics. Crucially, strictly preserving the original text indices ensures hallucination-free hallucination.
    • Edges ($\mathcal{E}$): represent the discourse linkages and dependency relations between EDUs (e.g., elaboration, contrast), capturing the logical flow often lost in standard retrieval.
  2. 2.
    Phase II: Sub-tree Retrieval and Linearization. A lightweight ranking module evaluates the relevance between the query $q$ and the structural nodes in $\mathcal{T}$. Instead of retrieving isolated sentences, it identifies the optimal task-relevant sub-trees $\mathcal{S}\subset\mathcal{T}$. Finally, these selected sub-trees are linearized back into a coherent text sequence $\mathcal{D}^{\prime}$. This results in a compressed context where $|\mathcal{D}^{\prime}|\ll|\mathcal{D}|$, retaining both global structural integrity and fine-grained details essential for downstream reasoning.
2.2 LingoEDU
We frame the task of Document Structure Analysis as a traceable context compression problem. As illustrated in Figure 2, our framework operates by transforming a linear discourse sequence into a condensed hierarchical tree, where every node is strictly anchored to the source via coordinate pointers.
Figure 2: Overview of the LingoEDU. (a) Coordinate System Construction: The continuous input text is segmented into EDUs, creating an addressable sequence where each unit carries a unique coordinate ID. (b) Traceable Generation: Unlike standard summarization, the model outputs Augmented Markdown. It performs compression by generating closed index intervals (e.g., $12\text{--}15]$) rather than regenerating body text, effectively indexing the content. (c) Tree Realization: The output is parsed into a hierarchical semantic tree $\mathcal{T}$. The explicit ID spans serve as unhallucinated anchors, ensuring the abstractive nodes remain strictly faithful to the source context.
2.2.1 Problem Formulation: Coordinate-Based Discourse Representation
To enable this coordinate-based operation (Figure 2(a)), we first segment the input document $\mathcal{D}$ into a sequence of atomic building blocks termed Elementary Discourse Units (EDUs). Formally, the document is represented as a sequence $\mathcal{U}=\{e_{1},e_{2},\dots,e_{N}\}$, where each unit $e_{i}$ constitutes a triplet:
$e_{i}=\left(t_{i},\mathtt{pos}_{i},\mathtt{id}_{i}\right)$ (1)
Here, $t_{i}$ represents the textual content (typically a coherent clause or sentence), $\mathtt{pos}_{i}$ denotes its physical grounding (character offsets), and crucially, $\mathtt{id}_{i}$ acts as a unique sequential index. This index establishes a Coordinate System, allowing the system to reference content by pointers. The core objective is to learn a decomposition function $f:\mathcal{E}\rightarrow\mathcal{T}$ that maps the linear EDU sequence to a hierarchical tree $\mathcal{T}$ (Figure 2(c)). Each node $n_{j}\in\mathcal{T}$ acts as a compressed semantic capsule defined as:
$n_{j}=\left(h_{j},l_{j},\sigma_{j}\right),\quad\text{where }\sigma_{j}=[\mathtt{id}_{\text{start}},\mathtt{id}_{\text{end}}]$ (2)
In this tuple, $h_{j}$ is the semantic abstract (e.g., a section title or summary), $l_{j}$ denotes the hierarchical depth, and $\sigma_{j}$ represents the EDU Span—a closed interval explicitly pointing to the source range in $\mathcal{E}$. By enforcing the constraint $\sigma_{j}\subseteq[1,N]$, we ensure referential integrity: the generated structure is a lossless index purely derived from the input context, effectively eliminating generative hallucinations.
2.2.2 Training Strategy and Data Synthesis
Since high-quality, fine-grained hierarchical annotations for long contexts are scarce, we introduce a scalable automated pipeline to synthesize training data. This pipeline leverages a strong LLM to distill the logic of summarization-and-indexing.
Bi-Level Task Decomposition.
Recognizing that faithfulness requires different cognitive capabilities for different structural types, we decouple the data generation into two distinct sub-tasks to prevent “instruction conflict” (conflating visual layout with semantic reasoning):
  1. 1.
    Explicit Layout Extraction: The model extracts objective structural cues (e.g., Markdown headers, HTML tags) to form the document skeleton. This task enables high certainty with low ambiguity.
  2. 2.
    Deep Semantic Segmentation: For large text blocks lacking explicit formatting, the model focuses purely on semantic shifts to delineate finer-grained functional sections. This requires deep reasoning to resolve high ambiguity.
The Solver-Critic Refinement Loop.
To ensure the synthesized labels are high-quality, we implement an Iterative Refinement Mechanism:
  • The Solver: Proposes a hierarchical decomposition, attempting to abstract detailed content into high-level semantic nodes.
  • The Critic: Audits the proposal specifically checking whether the generated abstract (Title) accurately reflects the assigned span $\sigma_{j}$ without hallucination or semantic drift.
This adversarial collaboration ensures that despite the high compression rate, the structural integrity of the training data remains intact.
2.2.3 Traceable Generation
We train an open-source LLM on this synthesized corpus. To further guarantee robustness during inference, we employ an Augmented Markdown Schema (Figure 2(b)). The model is trained to generate nodes in the following format:
$\text{Output}=\underbrace{\text{\#\#}}_{\text{Level}}\ \underbrace{[\text{id}_{\text{start}}\text{--}\text{id}_{\text{end}}]}_{\text{Traceable Anchor}}\ \underbrace{\text{Concept Title}}_{\text{Semantic Abstract}}$ (3)
This design achieves dual goals: (1) Token Efficiency, as long text blocks spanning multiple EDUs are compressed into minimal tokens; and (2) Hallucination Elimination, as the strict span format forces the model to rely on the coordinate system rather than free-form generation. Constraints are further enforced during decoding to ensure only valid numerical indices from $\mathcal{E}$ are generated.
2.3 Ranking Module
Once the document is decomposed into the structural tree $\mathcal{T}$, the challenge shifts to identifying which nodes contribute most to answering the user query $q$. We introduce a Budget-Aware Semantic Filter that leverages the Decomposer’s output to perform precise context selection.
2.3.1 Why Node-Level Ranking?
Instead of retrieving at the sentence level—which often results in context fragmentation—or the document level—which introduces excessive noise—we perform retrieval at the Node Level. Our nodes $n_{j}$ encapsulate semantic completeness via their spans, allowing the model to judge the relevance of larger logical blocks without processing the full text.
2.3.2 Plug-and-Play Relevance Scoring
We define a relevance scoring function $\phi(q,n_{j})$ to quantify the pertinence of each node. While our framework allows $\phi$ to be instantiated by state-of-the-art LLMs (e.g., GPT-4) via prompt-based scoring, such approaches are computationally prohibitive for scanning dense tree structures. To balance performance with efficiency, we employ a lightweight ranking model as a cost-effective surrogate. Specifically, we utilize an open-source model [^1] https://huggingface.co/Qwen/Qwen3-Reranker-0.6B to compute the relevance score $s_{j}$:
$s_{j}=\phi_{\theta}\left(q,\;h_{j}\oplus t_{\text{rep}}\right)$ (4)
where $h_{j}$ is the generated abstract (title) and $t_{\text{rep}}$ is the representative text snippet from the span $\sigma_{j}$. By using a compact model (e.g., 0.6B parameters) rather than large-scale LLM APIs, we achieve high-throughput filtering with negligible latency.
2.3.3 Budget-Aware Greedy Selection
To address the limitations of fixed Top-$K$ retrieval (which neglects token consumption), we employ a dynamic selection strategy bounded by a context budget $B_{\text{max}}$.
We sort all nodes in $\mathcal{T}$ by their score $s_{j}$ in descending order and select nodes into a candidate set $\mathcal{C}$:
$\mathcal{C}=\left\{n_{j}\mid\sum_{n\in\mathcal{C}}\text{Len}(\text{Retrieve}(\sigma_{n}))\leq B_{\text{max}}\right\}$ (5)
where $\text{Retrieve}(\sigma_{n})$ fetches the original text EDUs corresponding to the span $\mathtt{id}_{\text{start}},\mathtt{id}_{\text{end}}]$. This greedy strategy aligns the retrieved context density with the LLM’s optimal window size.
2.3.4 Linearization
A critical failure mode in standard RAG is the loss of discourse coherence when disjoint chunks are concatenated. Thanks to the explicit coordinates $\mathtt{id}$ provided by our Decomposer, we apply a Re-ordering Protocol. The selected spans in $\mathcal{C}$ are sorted by their original start indices $\mathtt{id}_{\text{start}}$ before concatenation. This restoration of logical order enables the downstream LLM to perform effective reasoning across discontinuous but structurally organized segments.
3 Experiments
3.1 Evaluation of Structural Integrity and Compression
In this section, we address the first research question: Can state-of-the-art LLMs effectively compress long contexts while preserving original structural information? We compare our proposed method against frontier LLMs and commercial parsing APIs on a newly constructed benchmark.
Benchmark Construction.
The absence of public benchmarks for fine-grained document structure analysis motivated us to construct a specific dataset named StructBench. We compiled a test set of 248 documents, covering diverse formats (Web pages, PDFs), languages (Chinese, English), and genres (e.g., government files, institutional reports, academic papers, and technical tutorials). Document lengths vary significantly, ranging from 300 to 50,000 words. To ensure high-quality ground truth, documents were parsed, sentence-segmented, and manually annotated for discourse structure by human experts. To enable fair comparison with baselines that may struggle with leaf-level details, we extracted the structural backbone (top-level hierarchy) from the annotations to serve as the ground truth. [^2] The dataset is available at https://huggingface.co/datasets/deeplang-ai/StructBench
Evaluation Metrics.
We employ two complementary metrics to evaluate structural fidelity:
  • Tree Edit Distance (TED) (ted_zhang1989simple): A micro-level metric that measures structural dissimilarity by computing the minimum number of edit operations (insertion, deletion, substitution) required to transform the predicted tree into the ground truth. Lower TED indicates more precise structural alignment.
  • Document Level Accuracy (DLA): A macro-level metric defined as $\mathrm{DLA}=\frac{\lvert D_{\text{match}}\rvert}{\lvert D_{\text{all}}\rvert}$, where $D_{\text{match}}$ represents the count of documents where the decomposed structure (backbone) perfectly matches the ground truth. This is a rigorous metric requiring zero structural errors.
Baselines.
We compare our LingoEDU (built on Qwen3-4B) against two categories of strong baselines. For all LLMs, we designed specific prompts to instruct them to output hierarchical JSON/Markdown structures:
  • Frontier LLMs: We evaluated SOTA models including GPT-4o/4.1 (gpt-4o), OpenAI o3/o4-mini (o3), Claude 3.7 Sonnet/4 (Claude4), DeepSeek-V3/R1 (deepseekai2025deepseekv3technicalreport), and Qwen3-235B (qwen3). All LLM results are averaged over three distinct runs.
  • Commercial Parsing APIs: We selected Jina Reader [^3] https://jina.ai/reader and Firecrawl [^4] https://www.firecrawl.dev, which are widely used for web-to-markdown conversion. We deployed test documents on a static server to allow URL-based access.
Implementation Details.
Our method utilizes Qwen3-4B as the backbone. The training involved a two-stage process: (1) Continued Pre-training on $\sim$100k synthetic samples to learn layout patterns, followed by (2) Supervised Fine-Tuning (SFT) on thousands of meticulously manually annotated documents to align with human intent. All experiments were conducted on a Linux operating system running on a high-performance server equipped with an Intel Xeon 2.3GHz CPU, 1960GB of memory, and 8 NVIDIA A800 GPUs, each with 80 GB of VRAM.
Table 1: Performance comparison on StructBench. $\ast$ indicates the model is accessed via API. $\dagger$ denotes local deployment for latency testing. Best results are bolded, and second-best are underlined. Our method achieves SOTA structural accuracy with significantly lower costs.
[…]
GPT-4o General LLM 8.53 36.29% 0.0210 -
GPT-4.1 9.14 37.90% 0.0168 -
OpenAI o3 8.01 35.48% 0.0168 -
OpenAI o4-mini 8.45 36.29% 0.0092 -
Claude-3.7-Sonnet 9.98 35.48% 0.0286 -
Claude-4 7.98 41.53% 0.0286 -
Gemini-2.5-flash 8.12 33.74% 0.0040 -
Gemini-2.5-pro 8.15 35.89% 0.0162 -
DeepSeek-V3 9.12 34.68% 0.0012 -
DeepSeek-R1 8.44 35.08% 0.0046 -
Qwen3-32B 8.55 34.01% 0.0012 10.17
Qwen3-235B 9.81 27.02% 0.0012 -
Jina-Reader Parser API 17.04 - 0.0004 -
Firecrawl 16.81 - 0.0007 -
Our Method (LingoEDU) Specialized 5.67 46.77% 0.0007 1.20
Analysis of Structural Integrity.
Table 1 highlights the superiority of our explicit training paradigm.
  • General LLMs struggle with precision: Even advanced reasoning models like OpenAI o3 and DeepSeek-R1 plateau around a TED of 8.0-9.0. Qualitative analysis reveals they often hallucinate non-existent sub-sections or flatten deep hierarchies to save generation tokens.
  • Parsing APIs lack semantic depth: While low-cost, Jina and Firecrawl exhibit high TED scores (>16), indicating they rely on shallow HTML tags and fail to capture implicit discourse structures found in complex PDFs.
  • Specialized Efficiency: Our LingoEDU Decomposer achieves a remarkable TED of 5.67 and DLA of 46.77%, significantly outperforming Claude-4 (+5.24% absolute DLA). This confirms that structural understanding requires dedicated supervision rather than emergent prompting.
Efficiency and Cost Analysis.
[…]
Ablation Studies.
Table 2 validates the effectiveness of our design. First, the performance gap between “Indices Only” and our method highlights that explicit text generation acts as a crucial semantic anchor. Second, the model exhibits strong data efficiency, maintaining over 94% of its peak performance even when scaled down to 20% of the training data.
Table 2: Ablation Studies. We analyze the impact of formulation and training data scale.
Ablation Setting Variant DLA (%)
Output Formulation Indices Only 33.47
Indices + Text (Ours) 46.77
Data Scaling 20% Data 44.35
50% Data 45.97
100% Data 46.77
3.2 Main Results on Downstream Long-Context Tasks
In this section, we address the second research question: Does structure-aware compression tangibly enhance performance for downstream tasks? We evaluate the utility of our EDU-based Context Compressor across two distinct scenarios: standard long-context benchmarks and complex Deep Search pipelines.
3.2.1 General Long-Context Understanding
We first evaluate our method on LongBench (bai-etal-2024-longbench), a comprehensive benchmark covering multi-document QA, summarization, and few-shot learning.
Baselines.
We compare three settings:
  1. 1.
    Standard: Feeds the full (or truncated) original context directly to the LLM.
  2. 2.
    Self-Sum: Uses the LLM itself to generate abstractive summaries of the context before processing.
  3. 3.
    Glyph(cheng2025glyph): Applies an implicit compression baseline (where applicable) serving as a reference for latent-based methods.
Analysis of Results.
Table 3 compares performance across Gemini-2.5-Pro and GPT-4.1 backbones. Overall, our EDU-based approach yields consistent improvements over the Standard baseline, with relative gains ($\Delta$) peaking at +23.66% on few-shot tasks.
  • Multi-Document QA (HotpotQA, Musique): Our method outperforms Self-Sum (e.g., +2.22 points on HotpotQA with GPT-4.1). Abstractive methods often lose critical entities required for multi-hop reasoning. In contrast, our explicit tree structure preserves precise evidence chains via original text indices, enabling the model to “look up” exact details without hallucination.
  • Summarization Tasks: While Self-Sum is naturally strong here, our method remains competitive (e.g., surpassing Self-Sum on MultiNews by +1.44 points). Notably, it significantly outperforms the external Glyph baseline, suggesting that retaining hierarchical structure is more effective than latent compression for information retention.
Table 3: Performance comparison on LongBench. Standard denotes the baseline without compression, Self-Sum utilizes LLM’s built-in summarization, and Ours (LingoEDU) represents our structure-aware compression. $\Delta$ indicates the relative improvement of Ours over the Standard baseline. Bold values indicate best performance; underlined values indicate second-best.
Task Type Dataset Glyph Gemini-2.5-Pro GPT-4.1
Standard Self-Sum Ours (LingoEDU) $\Delta$ Standard Self-Sum Ours (LingoEDU) $\Delta$
Multi-Doc QA HotpotQA 66.42 35.20 37.78 40.46 +14.94% 65.83 67.89 70.11 +6.50%
2WikiMQA 72.98 38.10 39.90 40.91 +7.38% 72.98 74.39 74.68 +2.33%
Musique - 28.55 30.77 31.22 +9.35% 51.90 53.48 54.86 +5.70%
DuReader - 7.15 7.79 8.12 +7.69% 21.80 23.51 25.34 +16.24%
Summarization GovReport 25.53 4.10 4.34 4.25 +2.44% 29.97 30.98 31.56 +2.94%
QMSum 19.78 15.80 16.53 16.17 +2.34% 22.84 22.53 23.30 +0.61%
MultiNews - 4.05 4.44 4.85 +19.75% 20.85 22.06 23.50 +5.80%
VCSum - 5.80 6.17 6.36 +9.66% 12.50 13.71 14.62 +8.96%
Few-shot TREC 82.62 46.50 49.00 57.50 +23.66% 77.00 80.50 80.00 +3.90%
TriviaQA 88.54 59.85 62.31 63.25 +1.25% 90.07 93.69 93.76 +4.10%
SAMSum - 20.45 21.89 23.80 +11.39% 39.20 40.79 41.68 +6.33%
LSHT - 26.10 29.50 35.48 +3.45% 48.60 50.50 52.50 +8.02%
3.2.2 Impact on Deep Search
Unlike standard retrieval, Deep Search involves aggregating information from diverse, often noisy web sources to answer complex, open-ended queries. We integrated the LingoEDU module into a Deep Search pipeline and evaluated it on two challenging benchmarks designed to push the limits of current LLMs:
  • High-Difficulty Reasoning (HLE): We utilize Humanity’s Last Exam (HLE) (phan2025humanity), an expert-curated benchmark designed to assess frontier-level academic competence. From the original set of 2,500 highly challenging questions spanning multiple disciplines, we focus on the subset of 2,154 text-only questions to evaluate deep reasoning capabilities.
  • Real-World Noise (BrowseComp-ZH): To test robustness in a messy information environment, we employ BrowseComp-ZH (zhou2025browsecomp). This is the first high-difficulty benchmark evaluating real-world web browsing and reasoning within the Chinese information ecosystem. It comprises 289 complex multi-hop queries across 11 domains (e.g., Film & TV, Technology, Medicine) often embedded in noisy web layouts.
Robustness to Web Noise.
Deep search engines frequently ingest raw web pages cluttered with ads, navigation bars, and irrelevant links. As shown in Table 4, on the noise-intensive BrowseComp-ZH, our method delivers dramatic gains. Specifically, Qwen3-235B and DeepSeek-V3.1 show relative improvements of 50.59% and 32.53%, respectively. This confirms that our Decomposer acts as a semantic filter: by identifying logical EDUs, it effectively prunes “structural dead branches” while preserving the core content, a capability that standard summarization lacks in such high-entropy Chinese web contexts.
Scaffolding for Complex Reasoning.
For the academic reasoning required by HLE, where answers depend on synthesizing multiple distant clues across disciplines, DeepSeek-R1 gains +51.11% relatively (from 9.0 to 13.6). This suggests that providing a cleaner, structure-aware context acts as a “reasoning scaffold.” It prevents the model from getting lost in irrelevant details (context drift), allowing it to focus its compute budget on logical deduction across valid evidence.
Compatibility with Frontier Models.
Critically, our method remains additive even for the most advanced models like GPT-5 and Gemini 3 Pro (Table 4, rows 4-6). This indicates that even as model capacity grows, handling unstructured noise and long-context reasoning remain bottlenecks. Our explicit compression provides a complementary signal that enhances the internal processing of SOTA LLMs.
Table 4: Ablation study of the LingoEDU module on Deep Search benchmarks. Accuracy scores (%) are reported. Base: Standard Deep Search without compression; Self-Sum: Query-focused summarization; Ours (LingoEDU): Structural decomposition. $\Delta$ denotes the relative improvement of Ours over the Base baseline. Bold indicates best performance per row.
Model Backbone HLE (Academic Reasoning) BrowseComp-ZH (Noisy Web)
Base Self-Sum Ours (LingoEDU) $\Delta$ Base Self-Sum Ours (LingoEDU) $\Delta$
DeepSeek-R1 9.0 9.5 13.6 +51.11% 18.8 19.4 20.4 +8.51%
Qwen3-235B-Thinking 14.2 14.7 15.5 +9.15% 8.5 9.0 12.8 +50.59%
DeepSeek-V3.1 14.5 14.8 15.6 +7.59% 29.2 29.7 38.7 +32.53%
Closed-Source Models
GPT-5 25.0 25.9 27.1 +8.40% 29.0 29.8 31.8 +9.66%
Claude Opus 4.1 14.0 14.8 15.5 +10.71% 20.8 21.5 23.2 +11.54%
Gemini 3 Pro 26.1 26.7 30.1 +15.33% 47.5 48.0 49.0 +3.16%
4 Related Work
4.1 Context Compression
As the input context length for LLMs grows, compressing long documents into efficient representations has become a critical challenge. Existing approaches can generally be categorized into semantic summarization, soft prompting, and token-level pruning.
Explicit Compression Methods.
Dominant approaches in this category operate on discrete tokens, filtering out low-informative content to reduce computational overhead. Early methods like Selective-Context li:2023selective and LLMLingua jiang2023:llmlingua utilize perplexity-based metrics to prune redundant tokens. Recent advances, such as LLMLingua-2 pan:2024llmlingua2 and TokenSkip xia2025tokenskip, move towards data distillation and controllable pruning for higher efficiency. While effective at reducing sequence length, these methods often operate on discrete tokens or rigid sentence boundaries, disrupting the local coherence of the text. Meanwhile, they typically focus on preserving the most important global information, while overlooking the original article’s structural information and fine-grained details.
Implicit Compression Methods.
Alternatively, implicit methods map long contexts into continuous vector spaces or latent states. Works like AutoCompressor (chevalier-etal-2023-adapting) and ICAE (ge2023context) compress text segments into soft prompts or memory slots. More extreme approaches, such as 500xCompressor (li2025500xcompressor) and Coconut hao2024training, push this further by performing reasoning directly in the latent space. However, recent studies (li2025admtreecompressinglengthycontext) show that implicit compression methods tend to have positional bias. This means they often ignore information from the beginning or middle of the context, focusing instead on the most noticeable content and overlooking less prominent details. Also, these implicit methods (cheng2025glyphscalingcontextwindows; wei2025deepseekocrcontextsopticalcompression) tend to lack flexibility, as they often require specially designed post-training processes or the use of latent vectors as new inputs. This limits the applicability of such techniques to advanced API-based models.
Unlike these methods, our design allows our method to preserve global structure while capturing fine-grained details, and crucially, since it operates without latent representations, it ensures seamless compatibility with API-based models.
4.2 Hallucinations in LLMs
Hallucination in LLMs remains a pervasive issue, characterized by the generation of non-factual or unfaithful content (hit-survey; si2025teaching). Recent research has shifted from simply viewing hallucination as a generation error to a more nuanced perspective of “controllable generation.” A substantial body of work has taxonomized the causes of hallucination si2025aligning; liu2025prosocial; jiSurveyHallucinationNatural2023; zhao-etal-2025-looking, treating hallucination mitigation as a debiasing task and striving to eliminate LLM hallucinations. However, strictly eliminating such uncertainty may compromise model creativity and usability. jiangSurveyLargeLanguage2024 propose an alternative view, regarding hallucination as a manifestation of creativity that requires control rather than simply eliminating it. In this way, to mitigate the hallucination, many works attempt to introduce external knowledge integration to establish controllable context for the generation process of LLMs, e.g., RAG technologies. However, these methods (zhangHallucinationMitigationRetrievalAugmented2025; fanSurveyRAGMeeting2024) often rely on embedding-based retrieval stage and struggle with noisy retrieval or context integration due to lack of the structural relations of the retrieved context. Different from these studies, we attempt to establish context for controllable generation through structured context compression. Thus, our structured context compression via the proposed EDU-based Context Compressor can preserve global foresight while capturing fine-grained details, ensuring less noisy information and reducing the hallucination on downstream tasks.
5 Conclusion
In this work, we present the EDU-based Context Compressor, a plug-and-play framework that bridges the gap between massive context windows and effective information utilization. By shifting from linear reduction to hierarchical, discourse-aware decomposition, our method explicitly filters noise while retaining precise evidence chains essential for complex reasoning. We validate our approach through the construction of StructBench, revealing that even state-of-the-art generalist LLMs struggle with fine-grained structural analysis compared to our specialized model. Extensive experiments confirm that our method not only surpasses existing compression baselines in standard retrieval tasks but also acts as a crucial “reasoning scaffold” for deep search agents, offering significant performance gains in noisy, real-world environments. Our findings suggest that explicit structural modeling is a key pathway to unlocking the full potential of LLMs in long-context scenarios. Future work will explore applying this structural paradigm to multi-modal contexts and dynamic agent memory management.
Appendix A Method Details
To provide a concrete understanding of our pipeline, we visualize the transformation process from linear text to a structured hierarchy in Figure 3.
Figure 3: Qualitative example of Traceable Context Compression. (a) Input Document. An example of a raw input document. (b) The Structured Document Tree. The model outputs a hierarchical index that reorganizes the linear EDUs into semantic clusters (Sections and Subsections). Note that the leaf nodes contain explicit coordinate pointers, allowing the system to faithfully retrieve the exact original text without hallucination.
Appendix B Prompt Template for LLM Baselines
To strictly pinpoint the structural understanding capabilities of general-purpose Large Language Models (LLMs), we utilized a unified zero-shot system prompt. This prompt essentially instructs the model to act as a parser, extracting the hierarchy without modifying the content.
The specific prompt content is presented in Box B.
Applicability Scope
It is important to clarify the distinct inference paradigms used in our experiments:
  • LLM-based Baselines (Applied): This prompt was applied to all general-purpose models listed in Table 1, including GPT-4o series, Claude series, DeepSeek series, and Qwen series.
  • Commercial APIs (Not Applied): Services like Jina Reader and Firecrawl operate as specialized black-box parsers. They ingest URLs or files and return structure via internal logic, rendering external prompting inapplicable.
  • EDU (Ours) (Not Applied): Unlike general-purpose LLMs that require detailed instructions (Prompt Engineering) to define the task, our model is explicitly trained via Supervised Fine-Tuning (SFT) for this specific objective. It accepts the raw document stream and outputs the structured tree end-to-end, relying on its internal parametric knowledge rather than prompt-based context.
System Prompt used for LLM Baselines Instruction: Work your way down through the article’s heading structure, outputting each level of heading in Markdown format. Use only the original headings—do not include body text, do not summarize, and do not rewrite. If a heading is split into multiple sentences due to punctuation, combine them into a single complete heading and output it as one entry. Output the heading structure top-down, carefully identifying hierarchical relationships and determining whether overly detailed levels are necessary to output. Return only the final result without any additional explanations.